OpenStudy (anonymous):

What is the center and radius of the following circle: x2 + (y – 6)2 = 50 A. C (6, 0), r = 7. 07 B. C (0, -6), r = 7. 07 C. C (0, 6), r = 50 D. C (0, 6), r = 7. 07

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 (h,k) is center r is radius (h,k)=(0,6) r=sqrt(50)

OpenStudy (anonymous):

so is it C?

OpenStudy (anonymous):

no, does sqrt(50)=50?

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

\[\sqrt{50}=50?\]

sam (.sam.):

D

OpenStudy (anonymous):

oh never mind it doesnt

OpenStudy (chriss):

\[\sqrt{50} \approx 7.07\]